. You need to have $50,000 at the end of 10 years. To accumulate this sum, you have decided to save a certain amount at the end of each of the next 10 years and deposit it in the bank. The bank pays 8 percent interest compounded annually for long-term deposits. How much will you have to save each year (to the nearest dollar)?


1
Expert's answer
2021-12-13T18:44:13-0500

"FV=\frac{R(1+r)^n-1}{r}=50000"


"R=\frac{FVr+1}{(1+r)^n}=\frac{50000\cdot0.08+1}{(1+0.08)^{10}}=\$1853.24\approx \$1853"